what formulas do i use for these problems:
Problem #5
John lee saving account has a balance of $2449 after 7 years what will the amount of interest be at 8% compounded semiannually?
Problem #7
Southwest dry cleaners believes that it will need new equipment in 5 years
the equipment will cost $26000 what lump sum should be invested today at 8% compounded semiannually to yield 26000?

Answer by stanbon(48551)

John lee saving account has a balance of $2449; after 7 years what will the amount of interest be at 8% compounded semiannually?
A(t) = P (1+r/n)^(nt)
A(7) = 2449(1+0.08/2)^(2*7)
A(7) = 4240.88
Interest earned = 4240.88 - 2449 = $1791.88
===============

Problem #7
Southwest dry cleaners believes that it will need new equipment in 5 years
the equipment will cost $26000 what lump sum should be invested today at 8% compounded semiannually to yield 26000?
---------
260000 = P(1+0.08/2)^(2*5)
P = $175646.68 (Amount to invest today)
===============
Cheers,
stan H.